A particular atom emits a photon of wavelength 471 nm when one of its electrons undergoes a transition from one stable energy level to another. If the final energy level has an energy of -6.04 eV, what is the energy of the inital energy level?


1
Expert's answer
2021-06-14T15:00:56-0400

Gives

Wavelength ("\lambda)" =471nm

Ef=-6.04ev

Ei=?

We know that

"E_f-E_i=h\upsilon"

"E_f-E_i=\frac{hc}{\lambda}"

"E_i=E_f-\frac{hc}{\lambda}"

Put value


"E_i=-6.04eV-\frac{6.625\times10^{-34}\times3\times10^8}{471\times10^{-9}\times1.6\times10^{-19}}"

"E_i=-6.04-2.64=-8.67eV"